# thanks: Corresponding authorthanks: left academia

### v1

@@ Abstract
We present a microscopic and dynamical model to study the formation, as well as dissociation and recombination processes of charmonium states in a heat bath at constant temperature and volume. Within this classical approach, the heavy quarks are described as Brownian particles in a background medium of light constituents and can therefore be described in terms of a Fokker-Planck equation with constant transport coefficients, which is then materialized through with relativistic Langevin simulations. The heavy quarks interact classically over a Coulomb-like screened potential to form a bound state if the relative energy of the pair becomes negative. Dissociation of bound states are possible as a result of screening effects of the potential as well as through scatterings with plasma particles. We demonstrate the full equilibration of the system and show that, for tuning the coupling constant, the obtained equilibrium charmonium yields are in accordance with the Statistical Hadronization Model.
@@ I Introduction
Pairs of heavy quarks and antiquarks are produced in primordial hard scatterings and their number is conserved throughout the whole medium evolution. Therefore information about their interaction history with the QGP can be obtained from measuring hadrons containing heavy quarks. Additionally, dissociation of quarkonium states may occur through elastic scatterings with medium particles and, at sufficiently high collision energies, these processes are possibly countered by regeneration of heavy quarks and antiquarks during the medium evolution Cao et al. (2013, 2015, 2016, 2018); Gossiaux et al. (2009); Nahrgang et al. (2014); He et al. (2012); van Hees et al. (2008); Plumari et al. (2012); Scardina et al. (2017); Cassing and Bratkovskaya (2008); Song et al. (2015) (for a summary of models and their comparison see Refs. Cao et al. (2019); Andronic et al. (2024)). A successful way to predict quarkonium yields measured in experiments at the hadronization phase boundary is the Statistical Hadronization Model (SHM) Andronic et al. (2003, 2018, 2021). This description neglects the dynamics that might occur during the medium evolution after the inital collision and instead focuses on describing the production of hadrons at the freeze-out stage of the QGP in the framework of statistical mechanics. This process is viewed as an instantaneous phenomenon governed by the chemical freeze-out temperature and the baryochemical potential, under the assumption of a statistically equilibrated source volume which can be described by a grand-canonical ensemble.
@@
In this work we propose a classical description of the heavy-quark motion, starting from a Fokker-Planck equation which is derived from the Boltzmann equation of the heavy-quark phase-space distribution and implemented through relativistic Langevin simulations. Within this approach the heavy quarks are approximated as Brownian particles and their motion results from random kicks due to elastic scatterings with light medium particles. The heavy quarks interact over a complex potential of which the real screened Coulomb-like part describes an attractive force between a heavy-quark-antiquark pair in order to form bound states, while the imaginary part takes interactions with the medium particles into account. In that way, the formation of charmonia as well as dissociation and recombination processes of bound states in the QGP can be described.
@@
The paper is organized as follows: In Sec. II the formalism of the Fokker-Planck equation is introduced and the realization of the heavy-quark motion in the form of a Langevin equation is described. Afterwards, Sec. III explains the potential between heavy quarks and antiquarks, which is able to form bound states. This (complex) potential can describe dissociation processes of bound states via its imaginary part, from which the drag coefficient is extracted. Section IV focuses on the formation of bound states, the binding energy of a charmonium state as well as the time evolution of the bound states under different conditions. Furthermore, the equilibration time of the charmonium states in connection with the drag coeffient is studied in Sec. V Finally, in Sec. VI the charmonium yields resulting from our model are compared to the predictions of the SHM. Finally, we present our conclusions and outlook in Sec. VII.
@@ II Fokker-Planck equation and Langevin simulations
The dynamics of heavy quarks in a thermal bath can be described by Brownian motion due to their large mass relative to the masses and the typical momentum exchanges with the light constituents of the QGP. The motion of heavy particles is viewed as random kicks resulting from momentum exchanges from elastic scatterings with light particles of the background medium. Under the assumption of local thermal equilibrium such an approach can be realized with a Fokker-Planck equation, describing the time evolution of the probability density of the heavy quarks under the influence of a drag and random forces. This equation can be derived from the relativistic Boltzmann equation for heavy-quark phase space distribution,
@@
In the limit of small exchanged momentum $\bm{k}$ as compared to $\bm{p}$, the collision integral can be expanded reducing the Boltzmann equation to a Fokker-Planck equation Svetitsky (1988); Pitaevskii and Lifshitz (2012); He et al. (2013a); Rapp and Hees (2010). Assuming, in addition, homogeneity and the absence of mean-field effects one gets,
@@
called the drag force and diffusion coefficients, which encode the forces acting on the heavy quark and its interactions with the medium. Imposing that the solution of the Fokker-Planck equation converges to the relativistic Boltzmann-Jüttner distribution in equilibrium, the two coefficients are connected via the dissipation-fluctuation theorem He et al. (2013b),
@@
where $E(p)=\sqrt{p^{2}+M_{c}^{2}}$ is the relativistic energy of the charm quark with mass $M_{c}$.
@@
Assuming these conditions the fluctuation-dissipation relation in (6) reduces to He et al. (2013b),
@@
In our following numerical approach, we consider a microscopical description of the system. This is realized through the relativistic Langevin equation van Hees et al. (2006) Dunkel and Hänggi (2009), and serves as the starting point for the practical simulation of heavy-quark propagation. The stochastic equations of motion for the heavy quarks are defined by the Langevin update steps, describing the change of position and momentum discretized in a time interval due to random momentum kicks from elastic scatterings with medium particles,
@@
Finally, we will find a simple relation between the friction from in the Langevin equation $\gamma(p)$ and the diffusion coefficient $D(p)$ by equating Eq. (21) to the fluctuation-dissipation relation (10). We obtain He et al. (2013b)
@@
While the coordinate update is straightforward, the momentum update requires a two-step computation due to the choice of the post-point scheme for the covariance matrix. Primarily, the momentum update in the pre-point scheme with $\zeta=0$ is computed, which can be utilized to calculate the momentum increment $dp_{j}^{\mathrm{drag}}$ with $\gamma$ and $D(p)$. Subsequently, the diffusion coefficient at $|\bm{p}+d\bm{p}|$ is evaluated in order to calculate $dp_{j}^{\mathrm{diffusion}}=\sqrt{2dtD(|\bm{p}+d\bm{p}|)}\ \rho_{j}$. Finally, the sum of $dp_{j}^{\mathrm{diffusion}}$ and $dp_{j}^{\mathrm{drag}}$ yields the total momentum increment for one full time step.
In the end Eq. (24) and (25) represent the well-known relativistic Langevin equation
@@
In the principle the white noise can or should be a coloured, non-Markovian noise with a non-local correlation in time Chen et al. (2023) Pooja et al. (2023). However, such general Langevin equations are much more numerically involved and thus for our present setting we will stay with the Markovian approximation, Eq. (26).
@@ III Heavy Quark Potential
Figure 1: Real part of the complex potential from Ref. Blaizot et al. (2016) as a function of the relative distance of the heavy-quark pair for different temperatures
@@
Figure 2: Upper panel: Drag coefficient as a function of temperature compared to results from the Texas A&M model Huggins and Rapp (2012) and the Catania quasiparticle model Scardina et al. (2017). Lower panel: Spatial diffusion coefficient of the charm quarks as a function of the temperature ($T_{c}=160\ \text{MeV}$) compared to the same models as well as data from lattice QCD Banerjee et al. (2012), Kaczmarek (2014) and pQCD calculations Moore and Teaney (2005), van Hees and Rapp (2005).
The left panel of Fig. 2 shows the drag coefficient $\gamma$ for different temperatures, illustrating an enhancement with increasing temperatures. Compared with other heavy-quark drag force determinations in the literature, our values are in the same physical range Beraudo et al. (2018). In this figure specifically we compare the drag coefficient resulting from the employed potential model with the results of the Texas A&M model Huggins and Rapp (2012) based on a $T$-matrix calculation using the heavy-quark internal energy from lattice-QCD as a potential, and the Catania quasiparticle model Scardina et al. (2017) extracted from simulations based on Langevin (LV) or Boltzmann (BM) equations. In the lower, right panel the spatial diffusion coefficient $D_{s}$ of the charm quarks is displayed, which is directly related to the mean-square displacement according to
@@
We note that the spatial diffusion coefficient in our simulation which results from the drag force proposed in Blaizot et al. (2016) appears to be smaller then in the models we are comparing to, but still lying within the uncertainties of the data from lattice QCD. The reason for the small magnitudes of $D_{s}$ is the application of a stronger coupling $\alpha_{s}$, as mentioned above.
@@
With this prescription one can now analyze in detail the microscopic formation, dissociation, the dynamical equilibration and the thermal population of charmonium states. We note that such a microscopic model has already been proposed originally by Shuryak and Young Young and Shuryak (2009),Young and Shuryak (2010) and also in Blaizot et al. (2016).
@@ IV Dynamical Bound State Formation
$E_{c\bar{c}}=E_{c}+E_{\bar{c}}+V(|\bm{r}_{c}-\bm{r}_{\bar{c}}|)-E_{\textrm{tot}}\ .$ (38)
This is checked at every time step, and for every pair combination. Once the system has equilibrated, the distribution of the relative energy is expected to be given by the classical density of states, multiplied with the Boltzmann factor in order to obtain the thermal distribution of states,
$\frac{dN_{c\bar{c}}}{dE_{c\bar{c}}}$ $=g(E_{c\bar{c}})\exp{\left(-\frac{E_{c\bar{c}}}{T}\right)}$ (39)
$=\int\frac{d^{3}rd^{3}p}{(2\pi\hbar)^{3}}\delta\left(E-\frac{p^{2}}{2\mu}-V(r)\right)\exp{\left(-\frac{E_{c\bar{c}}}{T}\right)}$ (40)
@@
V Momentum relaxation and equilibration time
As mentioned before, the drag force quantifies the amount of interactions of the heavy quarks with the surrounding medium. Therefore it determines the relaxation rate of the average particle’s momentum to its equilibrium value. The equilibration time of the system—to form the final charmonia states— should depend inversely on the drag coefficient in the Langevin equation and is of the same order as $\tau_{R}$. In order to study the influence of the drag coefficient on the equilibration time, different simulations with a single charm-anticharm pair are run with the parameter $\gamma$ scaled up by a factor $k=2,3$ and $5$. The results for the situation of one charm-quark pair are depicted in Fig. 8.
Figure 8: Time evolution of the fraction of bound states for different scalings of the drag coefficient $k\gamma$ within box simulations with a single charm-anticharm pair, at constant temperature $T=160\ \text{MeV}$ and volume $V=(10\ \text{fm})^{3}$.
From the analysis of the curves it can be concluded that a higher drag coefficient leads to a somewhat faster equilibration of the system. This might be expected since to achieve equilibration, particles need to relax their momentum first, which they do in time scales of the order of the inverse of the drag coefficient, $\tau_{R}=\gamma^{-1}$. However, full equilibration—understood as achieving a time-independent number of bound states—is dominated by a much slower process, the genuine two-body dynamics that makes the eventual generation of charmonia, and dominated by the time scales of the pairwise potential $V(r)$. Both effects, momentum relaxation and equilibration, are account in the Langevin equation. To quantify the latter, we make an exponential ansatz of the form,

### v3

@@ Abstract
We present a microscopic dynamical model to study the formation, dissociation, and recombination processes of charmonium states in a heat bath at constant temperature and volume. Within this classical approach, heavy quarks are described as Brownian particles in a background medium of light constituents and can therefore be modeled by a Fokker-Planck equation with constant transport coefficients, which is then implemented through relativistic Langevin simulations. The heavy quarks interact classically via a Coulomb-like screened potential to form a bound state if the relative energy of the pair becomes negative. Dissociation of bound states is possible as a result of screening effects on the potential as well as through scatterings with plasma particles. We demonstrate the full equilibration of the system and show that the resulting equilibrium charmonium yields are in accordance with the Statistical Hadronization Model.
@@ I Introduction
Pairs of heavy quarks and antiquarks are produced in primordial hard scatterings and their number is conserved throughout the whole medium evolution. Therefore information about their interaction history with the QGP can be obtained from measuring hadrons containing heavy quarks. Additionally, dissociation of quarkonium states may occur through quasi-elastic scatterings with medium particles and, at sufficiently high collision energies, these processes are possibly countered by regeneration of heavy quarks and antiquarks during the medium evolution Cao et al. (2013, 2015, 2016, 2018); Gossiaux et al. (2009); Nahrgang et al. (2014); He et al. (2012); van Hees et al. (2008); Du and Rapp (2022); Plumari et al. (2012); Scardina et al. (2017); Cassing and Bratkovskaya (2008); Song et al. (2015); Pooja et al. (2024) (for a summary of models and their comparison see Refs. Cao et al. (2019); Andronic et al. (2024)). A successful way to predict quarkonium yields measured in experiments at the hadronization phase boundary is the Statistical Hadronization Model (SHM) Andronic et al. (2003, 2018, 2021). This description neglects the dynamics that might occur during the medium evolution after the initial collision and instead focuses on describing the production of hadrons at the freeze-out stage of the QGP in the framework of statistical mechanics. This process is viewed as an instantaneous phenomenon governed by the chemical freeze-out temperature and the baryochemical potential, under the assumption of a statistically equilibrated source volume which can be described by a grand-canonical ensemble.
@@
In this work we propose a classical description of the heavy-quark motion, starting from a Fokker-Planck equation which is derived from the Boltzmann equation of the heavy-quark phase-space distribution and implemented through relativistic Langevin simulations. Within this approach the heavy quarks are approximated as Brownian particles and their motion results from random kicks due to quasi-elastic scatterings with light medium particles. The heavy quarks interact over a complex potential of which the real screened Coulomb-like part describes an attractive force between a heavy-quark-antiquark pair in order to form bound states, while the imaginary part takes interactions with the medium particles into account. In that way, the formation of charmonia as well as dissociation and recombination processes of bound states in the QGP can be described.
@@
The paper is organized as follows: In Sec. II the formalism of the Fokker-Planck equation is introduced and the realization of the heavy-quark motion in the form of a Langevin equation is described. Afterwards, Sec. III explains the potential between heavy quarks and antiquarks, which is able to form bound states. This (complex) potential can describe dissociation processes of bound states via its imaginary part, from which the drag coefficient is extracted. Section IV focuses on the formation of bound states, the binding energy of a charmonium state as well as the time evolution of the bound states under different conditions. Furthermore, the relaxation time of the charmonium states in connection with the drag coefficient is studied in Sec. V Finally, in Sec. VI the charmonium yields resulting from our model are compared to the predictions of the SHM. Finally, we present our conclusions and outlook in Sec. VII.
@@ II Fokker-Planck equation and Langevin simulations
The dynamics of heavy quarks in a thermal bath can be described by Brownian motion due to their large mass relative to the masses and the typical momentum exchanges with the light constituents of the QGP. The motion of heavy particles is viewed as random kicks resulting from momentum exchanges from quasi-elastic scatterings with light particles of the background medium. Under the assumption of local thermal equilibrium such an approach can be realized with a Fokker-Planck equation, describing the time evolution of the probability density of the heavy quarks under the influence of a drag and random forces. This equation can be derived from the relativistic Boltzmann equation for heavy-quark phase space distribution,
@@
In the limit of small exchanged momentum $\bm{k}$ as compared to $\bm{p}$, the collision integral can be expanded reducing the Boltzmann equation to a Fokker-Planck equation Svetitsky (1988); Pitaevskii and Lifshitz (2012); He et al. (2013); Rapp and Hees (2010). Assuming, in addition, homogeneity and the absence of mean-field effects one gets,
@@
called the drag force and diffusion coefficients, which encode the forces acting on the heavy quark and its interactions with the medium. Imposing that the solution of the Fokker-Planck equation converges to the relativistic Boltzmann-Jüttner distribution in equilibrium, the two coefficients are connected via the dissipation-fluctuation theorem He et al. (2013),
@@
where $E(p)=\sqrt{p^{2}+M_{c}^{2}}$ is the relativistic energy of the charm quark with mass $M_{c}$.This was first derived in Pitaevskii and Lifshitz (2012) for the non-relativistic case.
@@
Assuming these conditions the fluctuation-dissipation relation in (6) reduces to He et al. (2013),
@@
In our following numerical approach, we consider a microscopical description of the system. This is realized through the relativistic Langevin equation van Hees et al. (2006) Dunkel and Hänggi (2009), and serves as the starting point for the practical simulation of heavy-quark propagation. The stochastic equations of motion for the heavy quarks are defined by the Langevin update steps, describing the change of position and momentum discretized in a time interval due to random momentum kicks from quasi-elastic scatterings with medium particles,
@@
Finally, we will find a simple relation between the friction from in the Langevin equation $\gamma(p)$ and the diffusion coefficient $D(p)$ by equating Eq. (21) to the fluctuation-dissipation relation (10). We obtain He et al. (2013)
@@
While the coordinate update is straightforward, the momentum update requires a two-step computation due to the choice of the post-point scheme ($\zeta=1$) for the covariance matrix. In the first step, a preliminary momentum increment, $dp_{j}^{\mathrm{drag}}$, is computed using $\gamma$ and $D(p)$. Subsequently, the diffusion coefficient is recalculated at the updated momentum, $|\bm{p}+d\bm{p}|$, in order to evaluate $dp_{j}^{\mathrm{diffusion}}=\sqrt{2dtD(|\bm{p}+d\bm{p}|)}\ \rho_{j}$. Finally, the total momentum increment for one full time step is obtained as the sum of $dp_{j}^{\mathrm{diffusion}}$ and $dp_{j}^{\mathrm{drag}}$. He et al. (2013); Klimontovich (1994); Dunkel and Hänggi (2009)
In the end Eq. (25) represent the well-known relativistic Langevin equation,
@@
In principle, the white noise could be replaced by a colored, non-Markovian noise with a non-local correlation in time Chen et al. (2023); Pooja et al. (2023). However, such standard Langevin equations are much more numerically involved and thus for our present setting we will stay with the Markovian approximation, Eq. (26).
@@ III Heavy Quark Potential
Figure 1: Real part of the complex potential from Ref. Blaizot et al. (2016) as a function of the relative distance of the heavy-quark pair for different temperatures. Notice that the potential is qualitatively similar to the one in Ref. Blaizot et al. (2016), but since modify the strong-coupling $\alpha_{s}$ (as explained in Sec. VI), it leads to a deeper potential for all considered temperatures.
@@
Figure 2: Upper panel: Drag coefficient as a function of temperature compared to results from the Texas A&M model Huggins and Rapp (2012) and the Catania quasi-particle model Scardina et al. (2017). Lower panel: Spatial diffusion coefficient of the charm quarks as a function of the temperature ($T_{c}=160\ \text{MeV}$) compared to the same models and to results from the Dynamical Quasi-Particle Model Song et al. (2024), as well as data from lattice QCD Banerjee et al. (2012); Altenkort et al. (2024) and pQCD calculations Moore and Teaney (2005); van Hees and Rapp (2005).
The left panel of Fig. 2 shows the drag coefficient $\gamma$ for different temperatures, illustrating an enhancement with increasing temperatures. Compared with other heavy-quark drag force determinations in the literature, our values are in the same physical range Beraudo et al. (2018). In this figure specifically we compare the drag coefficient resulting from the employed potential model with the results of the Texas A&M model Huggins and Rapp (2012) based on a $T$-matrix calculation using the heavy-quark internal energy from lattice-QCD as a potential, and the Catania quasi-particle model Scardina et al. (2017) extracted from simulations based on Langevin (LV) or Boltzmann (BM) equations. In the lower, right panel the spatial diffusion coefficient $D_{s}$ of the charm quarks is displayed, which is directly related to the mean-square displacement according to
@@
The spatial diffusion coefficient in our simulation, which results from the drag force proposed in Blaizot et al. (2016), appears to be smaller than the models we are comparing to, while the most recent data from lattice QCD Altenkort et al. (2024) supports smaller values of $D_{s}$. Possible reasons for the small magnitudes of $D_{s}$ is the application of a stronger coupling $\alpha_{s}$, as mentioned above, as well as using a charm-quark mass of $m_{c}=1.8\ \text{GeV}$, as opposed to the pQCD calculations Moore and Teaney (2005); van Hees and Rapp (2005) which were obtained using $M_{c}=1.4\ \text{GeV}$ and $M_{c}=1.5\ \text{GeV}$ respectively.
We note that the results for $D_{s}$ from the Catania group are not calculated using the drag force (34) like in our case, but in an independent way using $D_{s}=T^{2}/D(0)$, with the diffusion coefficient in momentum space $D(p)$ of Eq. (20). The differences between the two methods are described in Ref. Das et al. (2016).
@@
With this prescription one can now analyze in detail the microscopic formation, dissociation, the dynamical equilibration and the thermal population of charmonium states. We note that such a microscopic model has already been proposed originally by Shuryak and Young Young and Shuryak (2009, 2010) and also in Blaizot et al. (2016).
@@ IV Dynamical Bound State Formation
$\begin{split}E_{c\,\bar{c}}=E_{c}+E_{\bar{c}}+V(|\bm{r}_{c}-\bm{r}_{\bar{c}}|)-E_{\text{tot}}\\ =\sqrt{m_{c}^{2}+\bm{p}_{c}^{2}}+\sqrt{m_{\bar{c}}^{2}+\bm{p}_{\bar{c}}^{2}}+V(|\bm{r}_{c}-\bm{r}_{\bar{c}}|)\\ +\sqrt{(m_{c}+m_{\bar{c}})^{2}+(\bm{p}_{c}+\bm{p}_{\bar{c}})^{2}}\,.\end{split}$ (38)
This is checked at every time step, and for every pair combination. Once the system has thermalized, the distribution of the relative energy is expected to be given by the classical density of states, multiplied with the Boltzmann factor in order to obtain the thermal distribution of states,
$\frac{dN_{c\bar{c}}}{dE_{c\bar{c}}}$ $=g(E_{c\bar{c}})\exp{\left(-\frac{E_{c\bar{c}}}{T}\right)}$
$=\int\frac{d^{3}rd^{3}p}{(2\pi\hbar)^{3}}\delta\left(E-\frac{p_{rel}^{2}}{2\mu}-V(r)\right)\exp{\left(-\frac{E_{c\bar{c}}}{T}\right)}$
@@
In order to confirm that Eq.(38) is a suitable criterion for identifying charmonium states, we consider the probability distribution of the relative distance of the charm-anticharm quark pairs. In Fig. 4 we compare the distribution of all pairs, regardless of their relative energy (blue line) with the case where only pairs with a negative relative energy are included (dashed orange line). The comparison reveals that in both cases, a narrow peak at small distances emerges, corresponding to the relative distance between a quark and antiquark within a bound state. At larger distances, the potential between the heavy quarks becomes negligible, leading to a distance distribution depending on the box size, which is represented by the second peak of the blue line. Since this second peak is absent when only considering pairs with a negative relative energy, we conclude that Eq. (38) is a suitable benchmark for only pairs with a reasonably small relative distance as bound states with within our model.
Figure 4: The probability distribution of the relative distance between a charm-anticharm quark pair, calculated with the same parameters as in Fig. 3. The blue line depicts the distribution for all pairs, regardless of their relative energy, the orange dashed line corresponds to bound states, i.e. pairs with a negative relative energy.
@@
V Momentum relaxation and chemical equilibration time
As mentioned before, the drag force quantifies the amount of interactions of the heavy quarks with the surrounding medium. Therefore it determines the relaxation rate of the average particle’s momentum to its equilibrium value. The chemical equilibration time of the system—to form the final charmonia states— should depend inversely on the drag coefficient in the Langevin equation and is of the same order as $\tau_{R}$. In order to study the influence of the drag coefficient on the equilibration time, different simulations with a single charm-anticharm pair are run with the parameter $\gamma$ scaled up by a factor $k=2,3$ and $5$. The results for the situation of one charm-quark pair are depicted in Fig. 9.
Figure 9: Time evolution of the fraction of bound states for different scaling of the drag coefficient $k\gamma$ within box simulations with a single charm-anticharm pair, at constant temperature $T=160\ \text{MeV}$ and volume $V=(10\ \text{fm})^{3}$.
From the analysis of the curves it can be concluded that a higher drag coefficient leads to a somewhat faster chemical equilibration of the system. This might be expected since to achieve equilibration, particles need to relax their momentum first, which they do in time scales of the order of the inverse of the drag coefficient, $\tau_{R}=\gamma^{-1}$. However, full chemical equilibration—understood as achieving a time-independent number of bound states—is dominated by a much slower process, the genuine two-body dynamics that makes the eventual generation of charmonia, and dominated by the time scales of the pairwise potential $V(r)$. Both effects, momentum relaxation and chemical equilibration, are account in the Langevin equation. To quantify the latter, we make an exponential ansatz of the form,